OpenStudy (homeworksucks):

A ship travelling 55* [W of N] is 65.0 km farther after 3.0 h. What is the ship's velocity?

OpenStudy (anonymous):

Velocity means speed and direction are wanted. speed is 65.0/3.0, direction is 55 o W of N?

OpenStudy (homeworksucks):

The answer isn't 6.01 m/s

OpenStudy (wolfe8):

Is it 21.7 km/h?

OpenStudy (homeworksucks):

38 km/h

OpenStudy (wolfe8):

This is mind-boggling for me. I would assume we need to do something with the direction but I can't figure it out right now. It's too late for me. I would suggest Google. Good luck.

OpenStudy (anonymous):

) If v is the speed of the ship (in km/h) its velocity is [ - v sin(55) , v cos(55) ]. What we are told means that (65 km) / (3,0 h) = v cos(55) Therefore, the speed is v = 65 / (3 cos(55)) = 37.77468... = 37.8 km/h The velocity is thus [ - 65 tan(55)/3 , 65/3 ] = [-30.9432... , 21.6666...] Properly rounded, that answer is [ -30.9 km/h , 21.7 km/h ] We may also express the answer as 37.8 km/h at 55 degrees [W of N]. http://ca.answers.yahoo.com/question/index?qid=20110318001811AA8Lfno

OpenStudy (anonymous):

Velocity must be 65 km/ 3 hr= 21.7 km/hr and in the direction the ship is traveling. SI, mks units would be m/s, if needed.

OpenStudy (wolfe8):

I guess the question wasn't clear about what distance that 65km is between.